OpenStudy (anonymous):

Im trying to figure out how to do this problem... I dont understand it, and would realy like to have someone show me how to do it... step by step. It is solve by completing the square- 4x^2-6x+5=0 any help would be greatly appreciated thanks.

OpenStudy (anonymous):

\begin{align} 4 x^2 - 6 x + 5 =0 \end{align} \begin{align} 4 x^2 - 6 x =-5 \end{align} \begin{align} 4 (x^2 - \frac{6}{4} x) =-5 \end{align} \begin{align} 4 (x^2 - \frac{3}{2} x) =-5 \end{align} So far, what I did make sense when we complete square , we half the term with 'x' and then square it \[-3/2 * 1/2=(3/4)^2\]---> 9/16 \begin{align} 4 (x^2 - \frac{3}{2} x+\frac{9}{16}) =-5 \end{align} add the same to other side \begin{align} 4 (x^2 - \frac{3}{2} x+\frac{9}{16}) =-5+4\frac{19}{6} \end{align} \begin{align} 4 (x - \frac{3}{4} )^2 =-5+4\frac{19}{6} \end{align}

OpenStudy (anonymous):

Does that all say math processing error to anyone else?

OpenStudy (radar):

Looks legal to me

OpenStudy (anonymous):

Well then. Carry on.

OpenStudy (radar):

Are you concerned about the (4) (19/6)??

OpenStudy (anonymous):

I sure as hell am... isnt it supposed to be (9/16)??

OpenStudy (anonymous):

and arent you suppoosed to add the -5+4... so the answer would be -1(9/16)??

OpenStudy (anonymous):

Let me rewrite last two line

OpenStudy (anonymous):

Now just clean it up

OpenStudy (radar):

that is 4*(9/16)

OpenStudy (anonymous):

\begin{align} 4 (x^2 - \frac{3}{2} x+\frac{9}{16}) =-5+4*\frac{9}{16} \end{align} \begin{align} 4 (x - \frac{3}{4} )^2 =-5+4*\frac{9}{16} \end{align}

OpenStudy (radar):

@bast0573 Pay particular to the 4 which is outside the paren on the left it applies to the 9/16 also and that is why it appears on the right.

OpenStudy (radar):

What you add to the left you add to the right.

OpenStudy (anonymous):

ok... so you would actually do the equationto solve it out?? which would come out to -5+(36/64)?